OpenStudy (anonymous):

how do solve this -----> -(5z+12)=18 ?

OpenStudy (across):

What have you tried so far?

OpenStudy (anonymous):

-(5z+12)=18 -5z-12=18 -5z=30 z=30/-5 z=-2

OpenStudy (anonymous):

thanks guys !!

OpenStudy (anonymous):

@gtvpa You are wrong because you have to distribute the negative sign in front of the parentheses first. (It's the same as having -1(5z+12))